# Generating Verifiable Chain of Thoughts from Exection-Traces

### v2

@@ 3.1.3. Rigorous Concept Deduplication and Quality Scoring
¡CoT¿ (1) Initialize distances: ’A’ to 0, all others to $\infty$. Priority queue starts with (’A’, 0). (2) Extract ’A’ from queue. Update distances for neighbors ’B’ and ’C’; add to queue. (3) Continue extracting minimum, updating neighbors until destination ’F’ reached. (4) Backtrack from predecessors to reconstruct path. Final cost is 7. ¡Predicted Output¿
@@
¡CoT¿ (1) The final path to ’F’ has cost 7. To deduce start, trace predecessors backwards. (2) The predecessor of ’F’ in cost-7 path must be ’E’. Edge cost from ’E’ to ’F’ is 2, so cost to reach ’E’ is 5. (3) Continue backtracking through ’C’, which confirms starting at ’A’ achieves total cost of 7. ¡Predicted Input¿
@@
Figure 3. Bi-directional CoT data format examples. Blue tags denote prompt components (instruction, function, question); green regions show model’s trace-grounded reasoning and prediction. Code appears in monospace, reasoning in normal text.
Function Signature   — Class Signature
@@
Figure 4. Signature format templates with extracted metadata. Generated signatures specify function/class names, parameter names and types, and return variable names with types (e.g., tree: tuple, codes: dict). This metadata is stored to validate code/test consistency and prevent hallucinated names or type mismatches.
Correct Format   — Prohibited Format
@@
Figure 5. Test format requirements. Correct format (left) enables clean I/O extraction for trace generation with direct function calls in assert statements, while prohibited format (right) complicates trace analysis with intermediate variable assignments.
3.1.4. Hierarchical Problem Synthesis
For each high-quality concept, we synthesize a complete problem artifact set through a carefully ordered five-step prompting pipeline. The ordering is designed to constrain the solution space at each step, preventing type and naming inconsistencies.
Step 1: Instruction Generation. Given a concept and its description, we prompt the teacher model to generate distinct natural language problem instructions. The prompt is engineered for diversity along three axes: (1) problem domain (mathematics, finance, data processing, algorithms), (2) computational approach (iterative, recursive, dynamic programming, functional), and (3) difficulty (”medium” targeting 30-60 LOC, ”hard” targeting 50-100+ LOC). The prompt includes explicit anti-overlap guidance to avoid paraphrased versions of the same problem.
Step 2: Signature Generation. Before generating any code or tests, we prompt the LLM to analyze the instruction and produce a formal signature skeleton specifying: (1) implementation type (standalone function vs. class), (2) function signature with parameter types and return type, or (3) class signature with constructor and methods. The signature prompt uses a strict formatting template enforced via few-shot examples, as shown in Figure 4. From each signature, we extract and store metadata including function names, parameter names and types, and crucially, return variable names with their types (e.g., encoding: dict[str, str]). This metadata serves as ground truth for validating consistency across code, tests, and generated CoT rationales. This signature-first design is deliberate: once input/output types and function names are fixed, subsequent code generation cannot hallucinate different types or names, eliminating import failures and ensuring test-code alignment.
Step 3: Code Generation. With the signature locked, we generate candidate solutions per instruction by prompting with both the instruction and signature. The prompt explicitly instructs: (1) strictly adhere to the signature, (2) vary computational approach across implementations, and (3) ensure completeness.
Step 4: Test Scenario Identification. We prompt the LLM to analyze the problem and produce a concise list of test scenarios (e.g., ”Test basic functionality”, ”Test empty input”). This improves test coverage by encouraging reasoning about requirements before writing code.
Step 5: Test Generation. We generate unit tests (3 test suites per instruction, each with up to 10 test functions) using a meticulously engineered prompt enforcing strict structural constraints. Each test is a top-level function (not a class method) containing exactly one assert statement with direct function calls using inline arguments—no variable assignments. Figure 5 illustrates the correct format versus prohibited patterns. These constraints are critical for Stage B granular pass/fail clustering and Stage C clean trace extraction, while enabling parallel sandboxed execution.
3.2. Stage B: Execution-Based Verification and Agreement Clustering
The hierarchical generation in Stage A, while carefully constrained, produces noise: some candidate solutions are incorrect (due to LLM generation errors), some tests are malformed or have wrong expected outputs, and some instruction-signature pairs are ambiguous. Before generating expensive CoT rationales in Stage C, we must identify and discard low-quality artifacts. We adapt the Dual Agreement verification methodology from CodeT (chen2022codetcodegenerationgenerated), originally designed for single-solution selection, into a scalable batch filtering algorithm.
3.2.1. Mass Execution and Pass/Fail Matrix Construction
For each problem (identified by task_id), we have $m$ candidate solutions and $n$ candidate tests. We execute all $m\times n$ solution-test pairs in a secure, sandboxed environment (Docker containers with resource limits) to construct a binary pass/fail matrix $M\in\{0,1\}^{m\times n}$, where $M[i,j]=1$ if solution $i$ passes test $j$, and 0 otherwise. Execution failures (timeouts, runtime errors, assertion failures) are all treated as $M[i,j]=0$. This brute-force execution is computationally expensive ($O(m\cdot n)$ per task) but embarrassingly parallel.
3.2.2. Dual Agreement Clustering
The core insight of Dual Agreement is: If a large set of independently generated solutions all pass the same large set of independently generated tests, it is statistically unlikely that both the solutions and tests are incorrect. We formalize this as follows:
Clustering by Test Agreement. We partition the $m$ candidate solutions into clusters $\{C_{1},C_{2},\ldots,C_{k}\}$ such that all solutions within a cluster $C_{i}$ have identical pass/fail patterns across all $n$ tests. Formally, solutions $s_{a}$ and $s_{b}$ are in the same cluster iff:
$\forall j\in[1,n]:M[a,j]=M[b,j]$
This is implemented via hash-based grouping: we compute a fingerprint for each solution’s test pass pattern (bit vector hash), then group solutions by fingerprint.
Cluster Scoring. For each cluster $C_{i}$, let $T_{p}(C_{i})$ denote the set of tests that all solutions in $C_{i}$ pass. We assign a quality score:
$\text{Score}(C_{i})=|C_{i}|\times|T_{p}(C_{i})|$
This score captures two dimensions: (1) Solution agreement ($|C_{i}|$): If many independently generated solutions have identical behavior, they likely implement the same correct algorithm. (2) Test coverage ($|T_{p}(C_{i})|$): If they pass many tests, the tests likely have correct expected outputs.
Best Cluster Selection. For each task, we rank clusters by score and select the single highest-scoring cluster. We extract one canonical solution from this cluster (selecting the pair with maximum code coverage) and its associated passing tests. Algorithm 1 formalizes this procedure. This method effectively filters multiple noise sources: incorrect solutions land in low-scoring clusters due to failing many tests (low $|T_{p}|$), malformed tests reduce scores but cannot eliminate correct clusters if other valid tests exist, and ambiguous instructions cause solutions to fragment into multiple small clusters (low $|C_{i}|$), with the highest-scoring cluster representing the most common interpretation.
Algorithm 1 Dual Agreement Clustering
0: Solutions $S=\{s_{1},\ldots,s_{m}\}$, Tests $T=\{t_{1},\ldots,t_{n}\}$
0: Verified solution $s^{*}$ and test suite $T^{*}$
1: Execute all $m\times n$ pairs to build matrix $M\in\{0,1\}^{m\times n}$
@@ 3.2.2. Dual Agreement Clustering
3:for each solution $s_{i}\in S$ do
4:  $\textit{pattern}_{i}\leftarrow M[i,:]$ // Test pass/fail pattern
5:  Add $s_{i}$ to cluster with matching pattern in clusters
6:end for
7:for each cluster $C_{i}\in\textit{clusters}$ do
8:  $T_{p}(C_{i})\leftarrow\{t_{j}:\forall s\in C_{i},M[s,j]=1\}$ // Commonly passed tests
9:  $\textit{Score}(C_{i})\leftarrow|C_{i}|\times|T_{p}(C_{i})|$
10:end for
[…]
14:return $(s^{*},T^{*})$
3.2.3. Theoretical Justification: Why Multiple Candidates Matter
The effectiveness of Dual Agreement relies critically on generating multiple diverse candidates. While a single solution-test pair provides no signal about correctness, multiple independently generated candidates enable statistical consensus detection through two complementary properties.
Unlikely Collision Property. The probability that $K$ independently incorrect solutions would all pass the same incorrect test suite decreases exponentially with $K$. Formally, if each incorrect solution has probability $p$ of coincidentally passing an incorrect test, then the probability that all $K$ incorrect solutions pass is $p^{K}$. For reasonable values (e.g., $p=0.3$, $K=5$), this yields $p^{K}=0.00243$, making such false consensus highly unlikely. In contrast, $K$ correct solutions deterministically pass correct tests. This means large $|C_{i}|$ (solution agreement) and large $|T_{p}(C_{i})|$ (test coverage) together provide strong correctness signals—clusters with both high $|C_{i}|$ and high $|T_{p}|$ survive the scoring threshold, filtering out isolated or weakly-validated artifacts.
We can formalize why this scoring creates clear separation between correct and incorrect clusters. Let $C_{correct}$ be the cluster of correct solutions and $C_{error}$ be any cluster of incorrect solutions. For the correct cluster:
$\mathbb{E}[\text{Score}(C_{correct})]=\mathbb{E}[|C_{correct}|]\times n_{valid}$
where $n_{valid}$ is the number of valid tests. For an incorrect cluster:
$\mathbb{E}[\text{Score}(C_{error})]\leq\mathbb{E}[|C_{error}|]\times n_{valid}\times(1-\delta)$
where $\delta>0$ represents the fraction of tests that expose the bug. When we have $K$ diverse candidate solutions with $m$ correct among them, the expected score ratio becomes:
$\frac{\mathbb{E}[\text{Score}(C_{correct})]}{\mathbb{E}[\text{Score}(C_{error})]}\geq\frac{m}{(K-m)(1-\delta)}$
As $K$ increases with sufficient diversity, $m$ (correct solutions) grows while incorrect solutions fragment into multiple small clusters with different error patterns, each with low $|C_{error}|$. This multiplicative scoring creates exponential separation, allowing correct clusters to dominate even with noisy candidates. However, this requires diversity: generating varied candidates via temperature sampling, prompt variation, and varied computational approaches in Stage A ensures that correct implementations converge while incorrect ones fragment. The effectiveness of this approach is empirically validated in Section 5.1.
3.3. Stage C: Execution-Grounded CoT Generation
The final stage transforms each verified solution-test pair into a rich, conversational training dataset where every reasoning step is grounded in the program’s actual execution. This is not a single-prompt process, but a meticulous five-step automated pipeline.
3.3.1. Step 1: Execution Trace Generation
For each verified solution and selected passing test cases, we generate a ground-truth execution trace capturing the program’s dynamic behavior. We use pysnooper(pysnooper), a lightweight Python tracing library, to instrument the target function (identified via signature metadata from Stage A).
Instrumentation Strategy. Rather than tracing the entire program (which includes test framework overhead), we selectively instrument only the function under test. This approach is robust against issues like recursion or complex interprocedural calls because the structural constraints imposed during Code Generation (Stage A)—which enforce a flattened logic without nested functions—guarantee that all core computation resides within the instrumented scope. For a function solution, we decorate it with @pysnooper.snoop(output=trace_file), then execute the test case in a sandboxed subprocess with a strict timeout (10 seconds). This produces a line-by-line trace file detailing: (1) Call events: Function entry with input arguments, (2) Line execution: Each line number executed in sequence, (3) Variable modifications: Every assignment with variable name, old value, and new value (capturing state transitions), and (4) Return events: Function exit with return value.
This trace is the immutable ground truth: a factual record of the program’s operational semantics—what values variables held at each point, how those values changed from statement to statement, and which execution paths were taken. It is not an LLM’s explanation or interpretation, but the actual runtime behavior.
3.3.2. Step 2: Trace Sanitization and Normalization
Raw pysnooper logs contain formatting artifacts unsuitable for model consumption: ANSI color codes, timestamp prefixes, file path prefixes, and redundant whitespace. A dedicated sanitization script strips these artifacts via regex-based cleaning, producing a clean, plain-text trace representation. This is critical because LLMs are sensitive to formatting: spurious tokens can degrade generation quality.
3.3.3. Step 3: Ground-Truth I/O Extraction and Question Generation
Before generating CoT rationales, we establish the precise input and output for each test case. For simple test functions like def test_basic(): assert solution([1, 2], 2) == [1], we use regex parsing to extract input = [1, 2], 2 and expected_output = [1]. For complex tests involving setup logic, we prompt an LLM to analyze the test code and extract the effective input/output pair. With ground-truth I/O established, we then prompt an LLM to generate a pair of natural language questions: (1) Forward question: ”Given the input [1, 2, 3] and target=2, what does the function return?”, and (2) Backward question: ”What input would cause the function to return [1]?” The prompt explicitly instructs natural phrasing to add linguistic diversity to the dataset.
3.3.4. Step 4: Trace-Grounded Chain-of-Thought Generation
This is the core of Stage C, where we use the sanitized trace as the factual basis for generating reasoning chains.
Forward CoT Generation: Narrating the Trace. We prompt an LLM with: (1) the instruction and function code, (2) the forward question, and (3) the sanitized execution trace. The prompt instructs: ”Narrate the execution trace step-by-step, explaining how the input is transformed into the output. Extract the <Predicted Output> you deduce from the trace.” Critically, the model is not generating an explanation from scratch; it is translating a factual trace into natural language. Each statement in the rationale can be verified against the trace: variable values mentioned must match the recorded runtime values, state transitions described must correspond to logged changes between statements, and the sequence of operations must follow the actual control flow path executed.
Backward CoT Generation: Deductive Reasoning from the Trace. For the backward direction, we prompt the model with the same trace but ask it to perform deductive reasoning: ”Given the final output state in the trace, explain how this output could only have been reached from the initial input. Extract the <Predicted Input> you derive.” This forces the model to reason backwards through the trace, explaining how the final state constrains the initial input.
Why This Trace-Grounding Eliminates Hallucination. Traditional self-instruct methods prompt an LLM to explain code without execution, leading to ”plausible hallucinations” where the explanation sounds correct but contains subtle factual errors (e.g., claiming a variable was incremented when it was decremented). By anchoring the rationale in a trace, we guarantee factual correctness: (1) Variable values are verifiable: Each value mentioned in the rationale is recorded at runtime and can be cross-referenced against the trace log. (2) State transitions are verifiable: Claims about how variables changed between statements (e.g., ”the counter increases from 0 to 1”) can be confirmed by comparing consecutive trace entries showing the variable’s old and new values. (3) Control flow is verifiable: Descriptions of which branches or loops executed can be validated against the trace’s record of which lines were executed and in what order. This ensures the entire causal chain—from initial state through semantic operations to final state—is factually grounded.
3.3.5. Step 5: Final Assembly into Conversational Format
The generated components (questions, CoTs, predicted I/O) are assembled into structured training examples. We create three versions of the dataset: (1) Forward-only: Each sample contains <Instruction>, <Function>, <Forward Question>, <CoT>, <Predicted Output>; (2) Backward-only: Each sample contains <Instruction>, <Function>, <Backward Question>, <CoT>, <Predicted Input>; (3) Bi-directional: Each sample contains both forward and backward turns in a single multi-turn conversation, teaching the model to reason in both directions.
4. Experimental Design
Our experiments are designed to rigorously quantify the benefits of our execution-grounded data and to determine the optimal configurations for fine-tuning.
4.1. Models and Benchmarks
Our model selection was deliberately designed to test our method across two open-source foundation models of a similar scale but with distinct pre-training philosophies. We selected granite-3.3-8b-base (granite3.3-base), an enterprise-grade architecture trained exclusively on permissively licensed data, which validates our approach for real-world applications where data provenance is critical. Alongside it, we chose Qwen2.5-Coder-7B (hui2024qwen2), a state-of-the-art, code-native specialist, to test if our data provides a reasoning signal capable of enhancing even expert models. The ability to improve both a safe generalist and a top specialist demonstrates the fundamental and broad applicability of our method. We evaluate all models on LiveCodeBench (Execution) (jain2024livecodebench) and CruxEval (Input/Output) (gu2024cruxeval), benchmarks specifically chosen for their focus on code execution reasoning. We finetune the base models for 10 epochs with a starting LR of 2e-6. More details about the training setup are included in the Appendix.
4.2. Data Curation and Preparation
The foundation of our experiments is a large-scale, high-quality dataset generated by our synthesis pipeline.
  1. (1)
    Concept Sourcing and Scaling: The pipeline began by ingesting a diverse set of permissively-licensed documents, including those from the ”Free Programming Books” source in the StarCoder2-documentation dataset (starcoder2-doc) and curated lists of programming books. These sources cover a wide range of topics, from core Python language features and data structures to advanced algorithms. This process yielded an initial set of approximately 15,000 deduplicated concepts. From this seed, we hierarchically synthesized a large pool of programming problems: for each concept, we generated multiple distinct instructions, and for each instruction, multiple candidate solutions and unit tests.
    4.2.1. Hyperparameter Selection for Data Generation
    To determine the optimal number of candidate solutions and unit tests per problem, we conducted systematic experiments with varying configurations. We evaluated different numbers of candidate solutions (5, 10, 15, 20) and unit test cases (5, 10, 20, 30, 40, 50) per instruction. Based on our ablation studies balancing quality and computational cost, we selected 5 candidate solutions and 30 unit tests per problem. The 30 unit tests were generated in 3 sets of 10 unit tests each, as described in Section 3.1, ensuring comprehensive coverage of edge cases and core functionality while maintaining diversity through multiple independent generation rounds.
  2. (2)
    Verification and Deduplication: This large, raw set of candidates was passed through our execution-based verification stage (Stage B), resulting in approximately 85,000 code-test pairs that were confirmed to be functionally correct. A final deduplication pass on similar code snippets and test logic yielded our initial master dataset.
    CoT Formatting: Each sample in the master set was formatted into three distinct reasoning structures. A forward CoT sample presents a question like Given input X, what is the output? followed by a trace-narrated rationale. A backward CoT sample asks "What input could produce output Y?" followed by a deductive, trace-grounded rationale. Figure 3 illustrates both formats showing the complete structure: instruction, function, question, step-by-step trace-grounded reasoning, and predicted output/input. Finally, a bi-directional sample combines these into a single, multi-turn conversational format.
  3. (3)
    Difficulty-Based Subsetting: From the master dataset of 85k code-test pairs, we created three distinct subsets for our ablation studies, based on different filtering strategies:
    • Full Set (54k): This is the complete, verified, and deduplicated dataset, representing the broadest collection of concepts.
    • Model Answerability Set (25k): This high-quality subset was created via a novel filtering strategy: we kept only the problems from the 54k deduplicated set that an LLM failed to solve correctly. This creates a targeted curriculum focused on the model’s specific weaknesses.
    • Content-Rated Difficulty Set (18k): This set was further refined from the 25k set by using an LLM to rate the conceptual difficulty of each problem, retaining only those rated as ”medium” or ”hard”.
This curation process provides two key dimensions for our experiments: the data curation strategy (54k vs. 25k vs. 18k) and the reasoning direction (forward, backward, or bi-directional).
4.3. Ablation Study Plan
Our experiments are designed as a sequential, top-down filter to efficiently identify the data configuration. While testing every combination of our data subsets and reasoning formats would be ideal, such a full factorial experiment is computationally infeasible. Our sequential approach is a practical and methodologically sound alternative that allows us to isolate the impact of our two primary contributions: difficulty-filtered subsets and CoT directionality. We operate on the reasonable assumption that the data subset demonstrating the highest quality for forward CoT will also be the most effective foundation for the other CoT formats. All experiments are conducted in a Supervised Fine-Tuning (SFT) setting.
  1. (1)
    Data Curation Ablation: We first aim to identify the optimal data subset. We fine-tune each base model on the 54k, 25k, and 18k datasets using only the forward CoT samples and compare their performance.
  2. (2)
    CoT Direction Ablation: Using the winning data subset from the previous stage, we then investigate the impact of the CoT format by fine-tuning models on the forward-only, backward-only, and the full bi-directional versions of that dataset.
  3. (3)
    Model Generalization: By performing this two-stage process on both Granite and Qwen-Coder, we validate the generalization of our findings across different model architectures.
5. Results and Analysis
We conduct a series of systematic ablation studies to evaluate the effectiveness of our data synthesis pipeline and determine the optimal training configurations. Our experimental design follows a sequential, top-down filtering approach: we first identify the best-performing data curation strategy, then use that dataset to determine the most effective reasoning direction. This entire process is performed on Granite-3.3-8b-base, and then repeated for Qwen2.5-Coder-7B to validate the generalizability of our findings. We evaluate all models on LiveCodeBench (Execution) and CruxEval (Output and Input prediction).
5.1. Validating Dual Agreement
Before presenting the main experimental results, we validate the effectiveness of our Dual Agreement verification approach (Stage B) through a controlled study on 5,000 programming problems with ground-truth correct solutions. Using the hyperparameter configurations described in Section 5.1.1, we measured both consensus scores and alignment with ground truth to quantify how the dual dimensions of solution agreement and test coverage affect verification reliability.
5.1.1. Consensus Score Scaling with Candidates
Figure 6 shows a heatmap of the percentage of problems achieving high consensus scores (Score $\tau$ threshold) as a function of the number of solutions (x-axis) and test cases (y-axis). The heatmap reveals a clear pattern: the bottom-right region (high solutions, high tests) shows consistently high consensus rates (70-90% of problems), while the upper-left region (few solutions, few tests) shows near-zero consensus rates. This validates our hypothesis that both multiple diverse solutions and comprehensive test suites are necessary for reliable verification. Notably, increasing solutions from 5 to 20 yields a 3x increase in problems with high consensus scores when test coverage is adequate ($\geq 20$ tests). This empirically confirms the theoretical exponential decay property: more candidates exponentially reduce the probability of false consensus.
Figure 6. Heatmap showing percentage of problems achieving high consensus scores as a function of number of candidate solutions (x-axis) and test cases (y-axis). The bottom-right region (many solutions, many tests) achieves 70-90% consensus, while the top-left (few solutions, few tests) shows near-zero consensus, validating that both dimensions are necessary for reliable verification.
Table 2. Comprehensive evaluation results across models, datasets, and training configurations on CruxEval and LiveCodeBench-Exec benchmarks. The experiments follow a top-down approach, with the winning configuration from one stage used in the next. Best results for each stage are highlighted in bold. And best result across the exp. stages are highlighted in grey.
Model CoT Direc. Data Subset LCB-Exec CruxEval Output CruxEval Input
@@ 5.1.1. Consensus Score Scaling with Candidates
Qwen2.5-Coder-7B Base (default) N/A 46.3 45.3 52.12 47.5 49
@@
Note: All Pass@k scores are reported as percentages. Green subscript values show improvement over the Base (Pre-trained) baseline. Abbreviations: Filt. = Filtered, perf. = performing, Fwd: Forward only CoT samples, Bwd: Backward only CoT samples, FT: Supervised fine-tuned.
5.2. Impact of Data Curation
Our first primary ablation investigates the impact of data quality and size by fine-tuning models on three forward-only data subsets: a complete 54k sample set, a higher-quality 25k set filtered for correctness, and a challenging 18k set filtered for difficulty.

### v3

@@ 3.3 Stage C: Execution-Grounded CoT Generation & Verification
<CoT> 1. Initialize distances: ’A’ to 0, all others to $\infty$. Priority queue starts with (’A’, 0). 2. Extract ’A’ from queue. Update distances for neighbors ’B’ and ’C’; add to queue. 3. Continue extracting minimum, updating neighbors until destination ’F’ reached. 4. Backtrack from predecessors to reconstruct path. Final cost is 7. <Predicted Output>
@@
<CoT> 1. The final path to ’F’ has cost 7. To deduce start, trace predecessors backwards. 2. The predecessor of ’F’ in cost-7 path must be ’E’. Edge cost from ’E’ to ’F’ is 2, so cost to reach ’E’ is 5. 3. Continue backtracking through ’C’, which confirms starting at ’A’ achieves total cost of 7. <Predicted Input>
@@
Figure 3: Forward and Backward CoT data format examples. Blue tags denote prompt components (instruction, function, question); green regions show model’s trace-grounded reasoning and prediction. Code appears in monospace, reasoning in normal text.
Function Signature   | Class Signature
@@
Figure 4: Signature format templates with extracted metadata. Generated signatures specify function/class names, parameter names and types, and return variable names with types (e.g., tree: tuple, codes: dict). This metadata is stored to validate code/test consistency and prevent hallucinated names or type mismatches.
Correct Format   | Prohibited Format
@@
Figure 5: Test format requirements. Correct format (left) enables clean I/O extraction for trace generation with direct function calls in assert statements, while prohibited format (right) complicates trace analysis with intermediate variable assignments.
Figure 6: Comparison against published baselines Rozière et al. (2024); Li et al. (2023); Guo et al. (2024); Wei et al. (2023b); Ding et al. (2024). Our results have hatch pattern. CL=CodeLlama, SC2=StarCoder2, DS=DeepSeekCoder, MCoder=MagicCoder, S=SemCoder, Py=Python, Inst=Instruct.
4 Experimental Setup
We use Mistral-Small-24B-Instruct Mistral AI (2025) for all data synthesis (code, tests, signatures, and CoT narration from traces). We fine-tune Granite-3.3-8B Mishra et al. (2024) and Qwen2.5-Coder-7B Hui et al. (2024) on our verified synthetic CoT data for 10 epochs with learning rate 2e-6, and report on LiveCodeBench-Exec (LCB) Jain et al. (2024) and CruxEval Gu et al. (2024) (Input/Output prediction) and HumanEval Chen et al. (2021). All results are reported as Pass@$k$ (fraction of problems solved within $k$ attempts); we use k=1k{=}1 unless otherwise stated. All evaluations use greedy decoding and default benchmark harness parameters.
Data Preparation. Our pipeline yielded 15,000 deduplicated concepts and 85,000 functionally correct code-test pairs after Stage B. Our verification mechanism (Value, Transition, and Control Flow Matching) then filtered these, rejecting approximately 30% of samples due to mismatches between generated CoTs and execution traces. Combined with deduplication and trace sanitization, this yielded 54,000 verified examples across three subsets: Full (54k), Model Answerability (25k) keeps only problems that the base model fails to solve before fine-tuning, removing trivial or already likely memorized samples, and Content-Rated Difficulty (18k) retaining medium/hard problems.
Contamination Control. We reduce benchmark overlap in three ways. Problems are synthesized from a concept curriculum derived from documentation and programming resources rather than benchmark prompts. The 25k Model Answerability subset described above further removes problems the base model already solves before fine-tuning. In addition, an LLM-based similarity screen on a representative random subset found no matches with LiveCodeBench or CruxEval under our overlap criteria. Together, these checks suggest that the synthesized training distribution is distinct from the evaluation sets.
@@ 4 Experimental Setup
Table 2: Comprehensive evaluation results (Pass@$k$, $k\in\{1,5\}$ attempts per problem). Best per stage in bold; best overall in grey. All Pass@$k$ scores are percentages; blue subscripts show improvement over Base. Fwd: Forward only CoT, Bwd: Backward only CoT, Bi-Directional: Forward + Backward, FT: Fine-tuned.
Model CoT Direc. Data Subset LCB-Exec CruxEval-O CruxEval-I
@@
Qwen2.5-Coder-7B Base (default) N/A 46.3 45.3 52.1 47.5 49.0
@@
5 Results
We structure our evaluation as a sequential ablation: we first compare the three curation subsets (54k, 25k, 18k) to identify the best-performing training set, then use it to evaluate forward, backward, and bi-directional reasoning, and finally ablate the verification approach itself. Table 2 presents all results.
5.1 Impact of Data Curation
We fine-tune models on three forward-only subsets to investigate data quality versus quantity.